1. Suppose today a 10 percent coupon bond sells at par. Two years from now, the required return on the same bond is 8 percent. What is the coupon rate on the bond? What is the YTM on the bond?

2. Vinya Inc. has 7.5 percent coupon bonds on the market that have 10 years left to maturity. The bonds make annual payments. If the YTM on these bonds is 8.75 percent, what is the current bond price? Assume the par value of a bond is $1,000. Do not round intermediate calculations and round your final answer to 2 decimal places.

3. Popie Drink Corp issued 12 year bonds 2 years ago at a coupon rate of 8.4 percent. The bonds make semi-annual payments. If these bonds currently sell for 105 percent of par value, what is the YTM? Assume the par value of a bond is $1,000. Do not round intermediate calculations and round your final answer to 2 decimal places.

4. Treasury bills are currently paying 7 percent and the inflation rate is 3.8 percent. What is the approximate real rate of interest? What is the exact real rate? Round your answer to 2 decimal places.
Answer #1

1)

Coupon rate remains the same.

Coupon rate = 10%

Since price sells at par, YTM will be equal to coupon.

YTM = 10%

2)

Coupon = 0.075 * 1000 = 75

Current bond price = Coupon * [1 - 1 / (1 + r)n] / r + FV / (1 + r)n

Current bond price = 75 * [1 - 1 / (1 + 0.0875)10] / 0.0875 + 1000 / (1 + 0.0875)10

Current bond price = 75 * 6.488886 + 432.222473

Current bond price = $918.89

3)

Number of periods = 10 * 2 = 20

Coupon = (0.084 * 1000) / 2 = 42

Price = 1.05 * 1000 = 1050

YTM = 7.67%

Keys to use in a financial calculator: 2nd I/Y 2, FV 1000, PV -1050, PMT 42, N 20, CPT I/Y